# NeuMaDiff: Neural Material Synthesis via Hyperdiffusion

### v1

@@ Abstract
High-quality material synthesis is essential for replicating complex surface properties to create realistic digital scenes. However, existing methods often suffer from inefficiencies in time and memory, require domain expertise, or demand extensive training data, with high-dimensional material data further constraining performance. Additionally, most approaches lack multi-modal guidance capabilities and standardized evaluation metrics, limiting control and comparability in synthesis tasks. To address these limitations, we propose NeuMaDiff, a novel neural material synthesis framework utilizing hyperdiffusion. Our method employs neural fields as a low-dimensional representation and incorporates a multi-modal conditional hyperdiffusion model to learn the distribution over material weights. This enables flexible guidance through inputs such as material type, text descriptions, or reference images, providing greater control over synthesis. To support future research, we contribute two new material datasets and introduce two BRDF distributional metrics for more rigorous evaluation. We demonstrate the effectiveness of NeuMaDiff through extensive experiments, including a novel statistics-based constrained synthesis approach, which enables the generation of materials of desired categories.
Figure 1: 3D models and scene rendered with our synthesized neural materials.
@@ 1 Introduction
Material synthesis plays a crucial role in visual computing, enabling the creation of realistic material appearances for applications in rendering, augmented reality, and scene understanding. Analytic models [42, 8, 53] are widely used to encode material appearance, due to their speed, memory efficiency, and ease of editing, which makes them suitable for real-time applications. However, their simplified representations often fail to capture the complexity of real-world materials, limiting their reconstruction accuracy. In contrast, measured data, while achieving unparalleled realism, is memory-intensive and challenging to manipulate, which restricts its flexibility. Implicit neural representations (INRs) offer a promising alternative, combining high-quality reconstructions with memory efficiency and a speed comparable with analytic models [49]. However, the lack of intuitive editing in INRs continues to hinder their usability and general adoption.
Deep-learning-based methods have also been explored and proven effective in generating highly realistic materials [6, 16, 64]. Nevertheless, they typically require extensive amounts of synthetic material data for training. Furthermore, their performance and efficiency are constrained by the high dimensionality of material data, which complicates the learning process. To overcome these challenges, our proposed pipeline, NeuMaDiff, introduces two strategies for effective data augmentation and employs implicit neural representations (INRs) as a low-dimensional, continuous representation of materials. This approach simplifies the learning process, enabling the model to capture the underlying material distribution more efficiently.
In addition, current material synthesis methods lack robust quantitative metrics for evaluating synthesis quality, making it difficult to assess and compare different approaches – unlike generative models in other fields [51, 4]. Another limitation is the absence of multi-modal conditioning, which would enable users to guide the synthesis process using diverse inputs, such as material type, text descriptions, or reference images. This limitation reduces the flexibility and control available to artists and designers, hindering the alignment of synthesized output with user intent. To address these limitations, we propose a set of novel BRDF distributional metrics and leverage a multi-modal conditional hyperdiffusion model to support flexible user input.
@@
  • A novel material synthesis pipeline using a multi-modal conditional hyperdiffusion model that supports user-specified material generation via material type, natural language descriptions, or image references;
  • Two new material datasets: AugMERL, an enhanced collection of tabulated BRDF values, and NeuMERL, a dataset of materials represented through implicit neural fields; and
  • A thorough evaluation of NeuMaDiff’s effectiveness, demonstrated through an extensive analysis that includes a set of novel BRDF distributional metrics and a novel constrained generation experiment to synthesize materials of desired categories.
@@ 2 Related Work
Material acquisition and databases
Capturing real-world material appearance requires lengthy acquisition times and substantial storage. Traditional capture uses four-axis gonioreflectometers [31, 58] and other system setups [55, 21, 10, 56]. Matusik et al. developed a high-resolution BRDF database with 100 isotropic materials [34]. Filip and Vávra contributed a database of 150 materials, many exhibiting anisotropic properties [15]. These material databases are continuously being augmented by later methods [38, 12]
@@
Material appearance has been widely represented by the bidirectional reflectance distribution function (BRDF) [39, 18, 36, 38, 57]. While analytical BRDF models [42, 8, 53, 5] offer efficient reconstruction and editing, their simplified assumptions limit the representation of complex real-world materials [38, 19]. Data-driven approaches offer higher realism [34], although they are often hard to manipulate and require large storage. Dimensionality reduction techniques can alleviate this issue, but at the expense of compromising material quality [29, 40]. Recently, deep learning methods provide efficient, low-dimensional BRDF representations with high-quality reconstruction [26, 61, 49, 17, 14, 20]. Our work leverages a neural field architecture [49] for efficient, realistic material modeling.
@@
Material synthesis based on analytic models has been widely explored [60, 35, 7, 6, 27, 50]. For data-driven representations, previous works have resourced to various strategies, including dimensionality reduction [34, 1], perceptual mappings [40, 46, 48] and deep learning [26, 17]: Hu et al. [26] used a convolutional autoencoder to learn a low-dimensional manifold from a measured BRDF database, enabling material synthesis and editing. However, their BRDF representation is constrained to a fixed resolution with high storage requirements. Similar to us, Gokbudak et al. [17] leveraged a hypernetwork architecture to predict the weights of a neural fields representation of material appearance. However, their approach requires sample measurements of BRDF data as input and is thus limited to sparse reconstruction. In this work, we introduce an approach for data-driven material generation based on a multi-modal hyperdiffusion architecture. Our method generates continuous and low-dimensional representations of material appearance, and can be conditioned on images, material categories and natural language descriptions.
@@ 3 Methods
An overview of our novel material synthesis pipeline, NeuMaDiff, is shown in Fig. 2, consisting of three main stages. First, we augment the MERL dataset through RGB permutation and PCA interpolation, generating the Augmented MERL (AugMERL) dataset (Sec. 3.1). Next, we adopt neural fields as a low-dimensional, continuous representation for materials, fitting them to individual materials in AugMERL to create a new dataset of neural material representations, Neural MERL (NeuMERL). Finally, we train a transformer-based, multi-modal hyperdiffusion model on NeuMERL to capture the complex distribution of neural materials, enabling high-fidelity and diverse synthesis through unconditional, multi-modal conditional, and constrained generation (Sec. 4).
Figure 2: An overview of NeuMaDiff, our novel neural material synthesis framework, consisting of three main stages. 1 (top left): Data augmentation using RGB permutation and PCA interpolation to create an expanded dataset, AugMERL; 2 (middle): Neural field fitted to individual materials, resulting in NeuMERL, a dataset of neural material representations; and 3 (bottom): Training a multi-modal conditional hyperdiffusion model on NeuMERL to enable conditional synthesis of high-quality, diverse materials guided by inputs such as material type, text descriptions, or reference images. We further propose a novel statistics-based constrained synthesis method to generate materials of a specified type (top right).
@@ 3.1 Data Augmentation
We utilize the MERL BRDF dataset [34], which includes 100 materials, each represented by densely sampled BRDF values in a $D_{\text{MERL}}=90\times 90\times 180$-dimensional space.
Through experimentation, we determined that 100 samples are insufficient for effective hyperdiffusion training. To address this, we augment the MERL dataset using RGB permutation and PCA interpolation. First, we permute the three color channels (RGB) of each MERL sample, yielding an expanded dataset of $100\times 3!=600$ samples. An example of RGB permutation is illustrated in Fig. 3.
@@
After applying RGB permutation, we perform principal component analysis (PCA) [1] to reduce the dimensionality of the BRDF data from $D_{\text{MERL}}$ to 300. In this lower-dimensional space, we perform linear interpolation to further augment the dataset, expanding it to $2400$ materials. Compared to direct linear interpolation in the high-dimensional BRDF space, interpolation in PCA space is more effective in capturing the underlying structure of the BRDF data and yields perceptually accurate results [34, 30, 44]. An example of materials generated through PCA interpolation is shown in Fig. 4. We refer to this augmented dataset as Augmented MERL (AugMERL). For additional details on PCA, please see Sec. A.1 in the supplementary.
@@ 3.2 Neural Field Fitting
Neural fields provide a low-dimensional, continuous representation for material data. Following prior work [49], we overfit a compact neural field $f_{r}^{{\bm{\mathbf{\xi}}}}$ parameterized by ${\bm{\mathbf{\xi}}}$, to each material in AugMERL. Once fitted, we treat the flattened weights of each neural field as the material’s neural representation. With 2400 materials in AugMERL, this process yields a dataset of 2400 neural material representations, which we refer to as Neural MERL (NeuMERL).
@@
In our preliminary experiments and other studies [e.g., 63], it is observed that directly using the conventional BRDF input format – namely the incident and outgoing directions ${\bm{\mathbf{\omega}}}_{i},{\bm{\mathbf{\omega}}}_{o}\in\mathbb{R}^{3}$ – can complicate the fitting process for certain materials and occasionally introduce undesirable artifacts possibly due to the high dimensionality of the input. To address this, we employ the Rusinkiewicz reparametrization [45], which defines the half and difference vectors ${\bm{\mathbf{h}}}$ and ${\bm{\mathbf{d}}}$ as follows:
${\bm{\mathbf{h}}}$ $:=\frac{{\bm{\mathbf{\omega}}}_{i}+{\bm{\mathbf{\omega}}}_{o}}{\lVert{\bm{\mathbf{\omega}}}_{i}+{\bm{\mathbf{\omega}}}_{o}\rVert};$ (1)
${\bm{\mathbf{d}}}$ $:=R_{\hat{{\bm{\mathbf{b}}}},-\theta_{\bm{\mathbf{h}}}}R_{\hat{{\bm{\mathbf{n}}}},-\varphi_{\bm{\mathbf{h}}}}{\bm{\mathbf{\omega}}}_{i},$ (2)
where $R_{{\bm{\mathbf{v}}},\alpha}$ denotes a rotation around the vector ${\bm{\mathbf{v}}}$ by the angle $\alpha$, $\hat{{\bm{\mathbf{n}}}}$ is the surface normal, and $\hat{{\bm{\mathbf{b}}}}$ is the surface binormal. This reparametrization helps improve the robustness of the neural field fitting process by addressing reciprocity constraints more directly [63].
@@
The high dynamic range of BRDF values makes fitting reflectance data particularly sensitive to error distribution. For low reflectance values, even minor fitting errors can have a large impact on the loss, causing shifts in perceived “hue” in rendered images, which leads to unrealistic colors and reduced visual fidelity. To address these issues, we employ a mean absolute logarithmic loss for BRDF values:
$\mathcal{L}_{\text{NF}}({\bm{\mathbf{\xi}}}):=\mathbb{E}_{\raisebox{-11.38109pt}{ $\mathclap{\begin{subarray}{c}\theta_{{\bm{\mathbf{h}}}},\theta_{{\bm{\mathbf{d}}}},\varphi_{{\bm{\mathbf{d}}}}\end{subarray}}$ }}\mkern-5.0mu\Big{[}\left|\log\left(1+f_{r}\cos\theta_{i}\right)-\log\left(1+f_{r}^{{\bm{\mathbf{\xi}}}}\cos\theta_{i}\right)\right|\Big{]}$
where $f_{r}$ denotes the ground-truth BRDF, which $f_{r}^{{\bm{\mathbf{\xi}}}}$ aims to approximate, and $\theta_{i}$ is the polar angle of the incident direction. This loss is computed per color channel, offering a balanced approach that stabilizes training across samples with both low and high values. Consequently, it enhances the model’s capability to manage dynamic reflectance variations, leading to more realistic color reproduction and improved visual fidelity.
@@
Ideally, the neural materials in NeuMERL should originate from a consistent distribution. However, due to a phenomenon known as weight symmetry [3], we observe that different weight configurations can yield the same neural field. For instance, swapping weights between two neurons in a hidden layer, or flipping the signs of both input and output weights for a neuron before an odd, linear, or piecewise-linear activation function like ReLU [37], results in an identical neural field. To address this, we propose using the optimized weights from the first fitted neural field as the initialization for subsequent neural field fittings. This approach helps align the weights across all fitted fields, promoting consistency within the NeuMERL dataset and facilitating smoother training of the hyperdiffusion model in the next stage.

### v2

# M3ashy: Multi-Modal Material Synthesis via Hyperdiffusion

@@ Abstract
High-quality material synthesis is essential for replicating complex surface properties to create realistic scenes. Despite advances in the generation of material appearance based on analytic models, the synthesis of real-world measured BRDFs remains largely unexplored. To address this challenge, we propose M3ashy, a novel multi-modal material synthesis framework based on hyperdiffusion. M3ashy enables high-quality reconstruction of complex real-world materials by leveraging neural fields as a compact continuous representation of BRDFs. Furthermore, our multi-modal conditional hyperdiffusion model allows for flexible material synthesis conditioned on material type, natural language descriptions, or reference images, providing greater user control over material generation. To support future research, we contribute two new material datasets and introduce two BRDF distributional metrics for more rigorous evaluation. We demonstrate the effectiveness of M3ashy through extensive experiments, including a novel statistics-based constrained synthesis, which enables the generation of materials of desired categories.
Figure 1: 3D models and scenes rendered with our synthesized neural materials demonstrate visually rich results.
@@ 1 Introduction
Material synthesis plays a crucial role in visual computing, enabling the creation of realistic material appearances for applications in scene understanding (gupta2013perceptual), material recognition (bell2015material), intrinsic image decomposition (bousseau2009intrinsic), generative image synthesis (karras2019style), and physics-based vision for simulation (wu2017marrnet). Material appearance is commonly modeled by the bidirectional reflectance distribution function (BRDF). Although there has been significant progress (e.g., gatys2015texture; zhou2018non) in generative modeling of analytic BRDFs (phong; ggx), there is a lack of work focusing on that of measured ones. For the analytic BRDFs, the actual per-point reflectance model is typically relatively simple and low-dimensional (ngan2005experimental; guarnera2016). In contrast, a measured BRDF is tabulated from real-world capture and can be substantially higher-dimensional (e.g., Matusik2003datadriven), with the ability to represent complex and irregular scattering behaviors that exceed the expressiveness of analytic models (ngan2005experimental). However, the high-dimensionality often hinders the performance of learning-based methods. To address this gap, we propose M3ashy, a framework for realistic material generation that leverages neural fields (sztrajman2021nbrdf) as an alternative low-dimensional continuous representation for material appearance that combines high-quality reconstruction with memory efficiency. This simplifies the learning process, enabling the model to capture the underlying material distribution more efficiently.
An additional challenge in material synthesis is the lack of robust quantitative metrics for evaluating synthesis quality, making it difficult to assess and compare different approaches, unlike generative models in other fields (theis2016evalgen; betzalel2022evalgen). A final limitation is the absence of multi-modal conditioning, which would enable users to guide the synthesis process using diverse inputs, such as material type, text descriptions or reference images. This limitation reduces the flexibility and control available to artists and designers. To address these limitations, we propose a set of novel BRDF distributional metrics and leverage a multi-modal conditional hyperdiffusion model to support flexible user input.
@@
  • A novel material synthesis pipeline using a multi-modal conditional hyperdiffusion model that supports user-specified material generation via material type, natural language descriptions, or image references.
  • A thorough evaluation of M3ashy’s effectiveness, including a set of novel BRDF distributional metrics and a novel constrained synthesis experiment to synthesize materials of desired categories.
  • Two new datasets: AugMERL, an enhanced collection of tabulated BRDF values, and NeuMERL, a dataset of materials represented through INRs.
@@ 2 Related Work
Material appearance has been widely modeled by the bidirectional reflectance distribution function (BRDF) (Nicodemus1977GeometricalCA; guarnera2016; montes2012overview; ngan2005experimental; westin2004comparison). While analytic BRDF models (phong; cook-torrance; ggx; disney) offer efficient reconstruction and editing, their simplified assumptions limit the representation of complex real-world materials (ngan2005experimental; guarnera2016; remapping2017; remapping2019). Data-driven approaches offer higher realism (Matusik2003datadriven), although they are often hard to manipulate and require large storage. Dimensionality reduction techniques can alleviate this issue, but at the expense of compromising material quality (lawrence2004efficient; NielsenPCA2015). Recently, deep learning methods provide efficient, low-dimensional representations (deepbrdf; zheng; sztrajman2021nbrdf; gokbudak2023hypernetworks; Fanlayered; metalayered). Our work leverages a neural field architecture (sztrajman2021nbrdf) for efficient, realistic material modeling. Please see Appendix˜A for additional related work on material acquisition and databases.
@@
Material synthesis based on analytic models has been widely explored (zhang2024textmat; memery2023materialNLP; Chen_2023_ICCV; Chen2023Text2TexTT; hu2023generating; tchapmi2022generating; xu23MATLABER; Henzler2021GenBRDFimg). For data-driven representations, previous works have resourced to various strategies, including dimensionality reduction (Matusik2003datadriven; abdi2010principal), perceptual mappings (NielsenPCA2015; serrano2016; sun2018connecting) and deep learning (deepbrdf; gokbudak2023hypernetworks), but these approaches do not offer a generative modeling of materials. deepbrdf use a convolutional autoencoder to learn a low-dimensional manifold from a measured BRDF database, enabling material editing. However, their BRDF representation is constrained to a fixed resolution with high storage requirements. gokbudak2023hypernetworks leverage a hypernetwork architecture to predict the weights of a neural fields representation of material appearance. Nevertheless, their approach requires sample measurements of BRDF data as input and is thus limited to sparse reconstruction. There also exist methods for text- (xu23MATLABER; memery2023materialNLP) and image-conditioned synthesis (deepbrdf; Henzler2021GenBRDFimg), but these either focus on analytic materials or do not provide a generative modeling approach.
In this work, we introduce a generative approach for measured real-world material based on a multi-modal hyperdiffusion architecture. Our method generates continuous, low-dimensional representations of materials, and can be conditioned on material type, natural language, and images. In Tab.˜1, we compare various state-of-the-art material modeling methods: Our approach is the first generative pipeline to support unconditional, multi-modal, and constrained synthesis of measured real-world materials while also contributing novel datasets and introducing quantitative metrics for material synthesis evaluation.
BRDF modeling method Measured Generative Type Text Image CS Datasets Metrics
DeepBRDF (deepbrdf)
Henzler2021GenBRDFimg
MATLABER (xu23MATLABER)
memery2023materialNLP
gokbudak2023hypernetworks
M3ashy (ours)
Table 1: Comparison of material modeling methods. Our M3ashy is the first generative pipeline for measured real-world materials that supports both unconditional and multi-modal conditional synthesis guided by type, text, or image. It also enables a statistics-based constrained synthesis (CS) and introduces novel datasets and material distributional metrics.
@@ 3 Methods
An overview of our material synthesis pipeline, M3ashy, is shown in Fig.˜2, consisting of three main stages. First, we augment the MERL dataset through RGB permutation and PCA interpolation, generating the Augmented MERL (AugMERL) dataset (Sec.˜3.1). Next, we adopt neural fields as a low-dimensional, continuous representation for materials, fitting them to individual materials in AugMERL to create a new dataset of neural material representations, Neural MERL (NeuMERL). Finally, we train a transformer-based, multi-modal hyperdiffusion model on NeuMERL to capture the complex distribution of neural materials, enabling high-fidelity and diverse synthesis through unconditional, multi-modal conditional, and constrained generation (Sec.˜4).
Figure 2: An overview of M3ashy, our novel neural material synthesis framework, consisting of three main stages. 1 (top left): Data augmentation using RGB permutation and PCA interpolation to create an expanded dataset, AugMERL; 2 (middle): Neural field fitted to individual materials, resulting in NeuMERL, a dataset of neural material representations; and 3 (bottom): Training a multi-modal conditional hyperdiffusion on NeuMERL to enable conditional synthesis of high-quality, diverse materials guided by inputs such as material type, text descriptions, or reference images. We further propose a novel statistics-based constrained synthesis method to generate materials of a specified type (top right).
@@ 3.1 Data Augmentation
We utilize the MERL dataset (Matusik2003datadriven), which includes 100 materials, each represented by $D_{\text{MERL}}=90\times 90\times 180$ densely sampled BRDF values.
Through experimentation, we determined that 100 samples are insufficient for effective hyperdiffusion training. To address this, we augment the MERL dataset using RGB permutation and PCA interpolation. First, we permute the three color channels (RGB) of each MERL sample, yielding an expanded dataset of $100\times 3!=600$ samples. An example of RGB permutation is illustrated in Fig.˜3.
@@
After applying RGB permutation, we perform principal component analysis (PCA) (abdi2010principal) to reduce the dimensionality of the BRDF data from $D_{\text{MERL}}$ to 300. In this lower-dimensional space, we perform linear interpolation to further augment the dataset, expanding it to $2400$ materials. Compared to direct linear interpolation in the high-dimensional BRDF space, interpolation in PCA space is more effective in capturing the underlying structure of the BRDF data and yields perceptually accurate results (Matusik2003datadriven; lawrence2004efficient; romeiro2010blind). An example of materials generated through PCA interpolation is shown in Fig.˜4. We refer to this augmented dataset as Augmented MERL (AugMERL). For additional details on PCA, please see Sec.˜B.1 in the supplementary.
@@ 3.2 Neural Field Fitting
Neural fields provide a low-dimensional, continuous representation for material data. Following prior work (sztrajman2021nbrdf), we overfit a compact neural field $f_{r}^{{\bm{\mathbf{\xi}}}}$ parameterized by $\bm{\mathbf{\xi}}$}, to each material in AugMERL. Once fitted, we treat the flattened weights of $f_{r}^{{\bm{\mathbf{\xi}}}}$ as the material’s neural representation. With 2400 materials in AugMERL, this process yields a dataset of 2400 neural material representations, which we refer to as Neural MERL (NeuMERL).
@@
In our preliminary experiments and other studies (e.g., sztrajman2021nbrdf; zhou2024physically), it is observed that directly using the conventional BRDF input format – namely the incident and outgoing directions ${\bm{\mathbf{\omega}}}_{i},{\bm{\mathbf{\omega}}}_{o}\in\mathbb{R}^{3}$ – can complicate the fitting process for certain materials and occasionally introduce undesirable artifacts possibly due to the high dimensionality of the input. To address this, we employ the Rusinkiewicz reparametrization (rusinkiewicz1998), which defines the half and difference vectors $\bm{\mathbf{h}}$} and $\bm{\mathbf{d}}$} as follows:
${\bm{\mathbf{h}}}\coloneqq\frac{{\bm{\mathbf{\omega}}}_{i}+{\bm{\mathbf{\omega}}}_{o}}{\lVert{\bm{\mathbf{\omega}}}_{i}+{\bm{\mathbf{\omega}}}_{o}\rVert};\,\,{\bm{\mathbf{d}}}\coloneqq R_{\hat{{\bm{\mathbf{b}}}},-\theta_{\bm{\mathbf{h}}}}R_{\hat{{\bm{\mathbf{n}}}},-\varphi_{\bm{\mathbf{h}}}}{\bm{\mathbf{\omega}}}_{i},$ (1)
where $R_{{\bm{\mathbf{v}}},\alpha}$ denotes a rotation around the vector $\bm{\mathbf{v}}$} by the angle $\alpha$, $\hat{{\bm{\mathbf{n}}}}$ is the surface normal, and $\hat{{\bm{\mathbf{b}}}}$ is the surface binormal. This reparametrization helps improve the robustness of the neural field fitting process by addressing reciprocity constraints more directly (sztrajman2021nbrdf; zhou2024physically).
@@
The high dynamic range of BRDF values makes fitting reflectance data particularly sensitive to error distribution. For low reflectance values, even minor fitting errors can have a large impact on the loss, causing shifts in perceived “hue” in rendered images, which leads to unrealistic colors and reduced visual fidelity. To address these issues, we employ a mean absolute logarithmic loss for BRDF values (sztrajman2021nbrdf):
$\mathcal{L}_{\text{NF}}({\bm{\mathbf{\xi}}})\coloneqq\mathbb{E}_{\raisebox{-11.38109pt}{ $\mathclap{\begin{subarray}{c}\theta_{{\bm{\mathbf{h}}}},\theta_{{\bm{\mathbf{d}}}},\varphi_{{\bm{\mathbf{d}}}}\end{subarray}}$ }}\mkern-5.0mu\Big[\left|\log\left(1+f_{r}\cos\theta_{i}\right)-\log\left(1+f_{r}^{{\bm{\mathbf{\xi}}}}\cos\theta_{i}\right)\right|\Big]$ (2)
where $f_{r}$ denotes the ground-truth BRDF, and $\theta_{i}$ is the polar angle of the incident direction. This loss is computed per color channel, offering a balanced approach that stabilizes training across samples with both low and high values. Consequently, it enhances the model’s capability to manage dynamic reflectance variations, leading to more realistic color reproduction and improved visual fidelity.
@@
Ideally, the neural materials in NeuMERL should originate from a consistent distribution. However, due to a phenomenon known as weight symmetry (liao2016important), we observe that different weights can yield the same neural field. For instance, swapping weights between two neurons in a hidden layer, or flipping the signs of both input and output weights for a neuron before an odd, linear, or piecewise-linear activation function like ReLU (nair2010rectified), results in an identical neural field. To address this, we propose using the optimized weights from the first fitted neural field as the initialization for subsequent neural field fittings. This approach helps align the weights across all fitted fields, promoting consistency within NeuMERL and facilitating smoother training of the hyperdiffusion model in the next stage.